OpenStudy (anonymous):

im stuck on what could equal 29? any ideas? 6a^2+29a-42 (2a+?)(3a-?)

OpenStudy (amistre64):

gotta be some factors of 49

OpenStudy (anonymous):

hmm like 7 and 7?

OpenStudy (amistre64):

7 and 6 maybe :)

OpenStudy (amistre64):

-42 ... i typoed it correctly

OpenStudy (anonymous):

:)ok bbut when i put them in the equation they don't equal 29

OpenStudy (amistre64):

you got the right front parts?

OpenStudy (anonymous):

lol i hope so

OpenStudy (anonymous):

ohh ohh mayb not

OpenStudy (anonymous):

if i factor out ? 2a

OpenStudy (anonymous):

I dk i get stuck in the break down and when i check it doesnt the middle number have to equal the product of the first and last?

OpenStudy (amistre64):

6*42 = 252 factors of 252 that subtract to 29 are: 1 252 ... 2 126 ... 4 63 ..... 3 84.... 6 42..... 12 21 .... 5 ..... 7 36; that looks promising -36 + 7 = -29 so (x-36/6) (x+7/6) (x-6) (x+7/6) (x-6) (6x+7)

OpenStudy (anonymous):

I sososososososo would have never come up with that?

OpenStudy (anonymous):

lets agree that this equation kinda sux

OpenStudy (anonymous):

grrrrrrrrrrrrrrr

OpenStudy (amistre64):

after al that work? nah, it was a glorious achievement; we should sacrifice 10000 virgins warriors in celebration of its completion lol

OpenStudy (anonymous):

-36 and possitive 7

OpenStudy (anonymous):

you mean, all you have to do is add up the numbers that equal 29? piece of cake.

OpenStudy (anonymous):

thanks amistre lol virgin warriors huh good luck on that achievement....;)))))

OpenStudy (anonymous):

oooh. sorry.

OpenStudy (anonymous):

no worries doglove

OpenStudy (anonymous):

that was too easy to be a question.

OpenStudy (anonymous):

hehe true but hey it was worth a try right!!

OpenStudy (anonymous):

yeah... i guess.

OpenStudy (anonymous):

ummm... what is the prime euuasion that equals 229?

OpenStudy (anonymous):

like this. exept this is for 20.\[5 \times 2 \times 2 =20.\]